OpenStudy (anonymous):

find vertical asymptote of f(x)= ((((x)^2)+(x)-(6)))/(((x)^2)-(x)-(2)))

OpenStudy (anonymous):

@satellite73 please help!

OpenStudy (anonymous):

can't read it, but whatever it is, set the denominator equal to zero and solve for \(x\)

OpenStudy (anonymous):

looks like the denominator is \(x^2-x-2\) put \[x^2-x-2=0\\ (x-2)(x+1)=0\] etc

OpenStudy (anonymous):

\[f(x) =\frac{ x^2 +x-6}{x^2-x-2 ? }\]

OpenStudy (anonymous):

yeah, ok, what i thought it might be i wrote most of the solution above you good from there?

OpenStudy (anonymous):

no.. here's what i got… when x=2 i get \[2^2+2-6 = 0\] when x=-1 i get \[(-1)^2 + (-1)-6 = -6 \neq 0\] correct?

OpenStudy (anonymous):

lets go slow

OpenStudy (anonymous):

for a rational function, the vertical asymptote is at the zeros of the denominator you cannot divide by zero

OpenStudy (anonymous):

the denominator of your rational function is \(x^2-x-2\) right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

therefore this question translates as "set \(x^2-x-2=0\) and solve for \(x\)"

OpenStudy (anonymous):

luckily for you, this one factors you can solve it by factoring as \[(x-2)(x+1)=0\] and that means \(x=2\) or \(x=-1\) those are both the zeros of the denominator, and also your vertical asymptotes

OpenStudy (anonymous):

okay so i don't have to take the limits?

OpenStudy (anonymous):

nope it just asks for vertical asymptotes the limits are either infinity or minus infinity, which are not numbers

OpenStudy (anonymous):

ok so i also have to get the H.A. for the same problem… can you tell me if i did it correctly?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

\[f(x) =\frac{ x^2 +x-6}{x^2-x-2 }\] just to refresh my memory

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty} \frac{ x^2+x-6 }{x^2-x-2} \frac{ \infty }{ \infty } form\]

OpenStudy (anonymous):

tmi degree of the numerator is the same as the degree of the denominator (they are both degree 2) so it is the ratio of the leading coefficients, which in this case is \(\frac{1}{1}=1\)

OpenStudy (anonymous):

so i took the highest power in denominator and after my workings by taking out limit i got

OpenStudy (anonymous):

i havent done that in class yet.. i was told i had to use highest power

OpenStudy (anonymous):

but i did get the answer to be 1 so perfect

OpenStudy (anonymous):

right the highest power of both is 2 since they are the same, take the ratio of the leading coeficients you can do it in your head for example the horizontal asymptote of \[\frac{4x^3+2x^2+x-1}{5x^3+20x-3}\] is \(y=\frac{4}{5}\)

OpenStudy (anonymous):

ohhhhh. ok.. makes sense now. sorry. so if i did out the limits for my V.A.s i should get infinity?

OpenStudy (anonymous):

oh yes either \(\infty\) or \(-\infty\)

OpenStudy (anonymous):

ok so i had to find va and ha for \[f(x) = \frac{ x+1 }{x^2-1}\] and i got va… x=1 and ha…. y=0….. is this correct?